# Generating 6-D Trajectories for Omnidirectional Multirotor Aerial Vehicles in Cluttered Environments

### v1

@@ I Introduction
In order to better exploit the potential of OMAVs, a proper trajectory generation framework is indispensable. It needs to generate 6-D $SE(3)$ trajectories and take into account the vehicle’s shape and pose to adapt to cluttered environment. Moreover, we expect this framework to have excellent computational efficiency.
@@
This paper proposes an efficient 3-stage 6-D trajectory generation framework for OMAVs in cluttered environments, as shown in Fig. 1. Similar to [11], 3-D polyhedral SFC is used to represent the collision-free region. First, we use existing methods to search for an initial feasible path from the start to the end point and generate a 3-D SFC based on the initial path. Then, the 6-D trajectory optimization stage efficiently generates a smooth, safe, and dynamically feasible 6-D trajectory connecting the start and the end states. For trajectory representation, we represent a vehicle’s attitude as a Hamilton unit quaternion $\mathbf{Q}\in\mathbb{S}^{3}$ [14] and parameterize it into a free 3-D vector using stereographic projection to eliminate the constraints inherent in the $\mathbb{S}^{3}$ manifold. This 3-D attitude vector is combined with the 3-D position vector to form a 6-D pose vector. The complete $SE(3)$ trajectory is then represented as a piece-wise polynomial of the pose vector over time. The vehicle’s shape and attitude are taken into account to apply whole-body safety constraints to the trajectory. We formulate the trajectory generation problem as a constrained optimization problem and finally transform it into an unconstrained one that can be solved using quasi-Newton methods.
@@
  • We propose a stereographic-projection-based method to represent rotation trajectories as curves in $\mathbb{R}^{3}$.
  • Based on the above trajectory representation, we propose an efficient 6-D trajectory optimization method that considers constraints including whole-body safety and dynamic limits for OMAVs.
  • We present an efficient 6-D trajectory generation framework that can give full play to the obstacle avoidance potential of OMAVs for the first time.
@@ II-B System Modeling of OMAVs
and we select CoM position $\mathbf{p}$, CoM velocity $\mathbf{v}=\dot{\mathbf{p}}$, $\mathcal{F}_{b}$’s orientation $\bm{\varepsilon}$ (expressed as Euler angles), and $\mathcal{F}_{b}$’s angular velocity $\bm{\omega}$ (all expressed in $\mathcal{F}_{W}$) as its state variables:
$\mathbf{x}:=[\underbrace{p_{x},p_{y},p_{z}}_{\mathbf{p}^{\top}},\underbrace{\phi,\theta,\psi}_{\bm{\varepsilon}^{\top}},\underbrace{v_{x},v_{y},v_{z}}_{\mathbf{v}^{\top}},\underbrace{\omega_{x},\omega_{y},\omega_{z}}_{\bm{\omega}^{\top}}]^{\top}\in\mathbb{R}^{12}.$ (2)
The system’s output is given by the position of its CoM and the orientation of $\mathcal{F}_{b}$ expressed in $\mathcal{F}_{W}$:
$\mathbf{y}(t):=\{\mathbf{p}(t),\mathbf{R}(t)\}\in SE(3).$ (3)
Expressing a rotation trajectory as a function of the rotation matrix w.r.t. time $t$ is intuitive. However, this implies equality constraints inherent in $SO(3)$, which can be troublesome for trajectory optimization. So, we consider parameterizing each rotation $\mathbf{R}$ as an unconstrained 3-D vector. Denote a rotation trajectory by a curve $\bm{\sigma}(t):[t_{0},t_{M}]\rightarrow\mathbb{R}^{3}$, and the corresponding $\mathbf{R}(t)$ is determined by a smooth surjection $\mathbf{R}(\bm{\sigma}):\mathbb{R}^{3}\rightarrow SO(3)$. Then, we express the 6-D trajectory as
@@
where $(\cdot)^{\wedge}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ denotes the skew-symmetric matrix of a 3-D vector and thus $\mathbf{a}\times\mathbf{b}=\mathbf{a}^{\wedge}\mathbf{b},\forall\mathbf{a},\mathbf{b}\in\mathbb{R}^{3}$. The inverse map is denote as $(\cdot)^{\vee}:\mathfrak{so}(3)\rightarrow\mathbb{R}^{3}$. The angular acceleration $\dot{\bm{\omega}}$ can be further calculated according to (5) :
@@
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \bm{\varepsilon}\\ \dot{\mathbf{p}}\\ \bm{\omega}\end{bmatrix}=\Psi_{x}(\mathbf{z},\dot{\mathbf{z}}):=\begin{bmatrix}\mathbf{p}\\ \bm{\varepsilon}(\bm{\sigma})\\ \dot{\mathbf{p}}\\ \left(\sum_{i=1}^{3}\frac{\partial\mathbf{R}(\bm{\sigma})}{\partial\sigma_{i}}\dot{\sigma}_{i}\mathbf{R}(\bm{\sigma})^{\top}\right)^{\vee}\end{bmatrix}.$ (7)
@@
Equ. (7) and Equ. (8) indicate that our selected trajectory representation $\mathbf{z}(t)$ has the properties as a flat output [15]. Therefore, it is convenient to transform constraints on the state $\mathbf{x}\in\mathbb{R}^{12}$ and the input $\mathbf{u}\in\mathbb{R}^{6}$ to constraints on $\mathbf{z}\in\mathbb{R}^{6}$ and $\mathbf{z}$’s finite derivatives using Equ. (7) and Equ. (8).
@@ II-C Position and Attitude Control
$\mathbf{v}_{\text{sp}}=\mathbf{K}_{P,p}\odot\bm{e}_{p},$ (9)
where $\odot$ represents the entry-wise multiplication of two vectors with the same dimensions. $\mathbf{K}_{P,p}\in\mathbb{R}^{3}$ is the proportional gains. The inner velocity control loop uses PID control with a feed-forward design, given as
$\mathbf{a}_{\text{sp}}=\mathbf{K}_{P,v}\odot\bm{e}_{v}+\mathbf{K}_{I,v}\odot\int\bm{e}_{v}\text{d}t+\mathbf{K}_{D,v}\odot\frac{\text{d}\bm{e}_{v}}{\text{d}t}+\mathbf{K}_{F,v}\odot\mathbf{v}_{\text{ref}},$ (10)
where $\mathbf{K}_{P,v},\mathbf{K}_{I,v},\mathbf{K}_{D,v},\mathbf{K}_{F,v}\in\mathbb{R}^{3}$ are gains for proportional, integral, derivative, and feed-forward control, respectively. $\bm{e}_{v}:=\bm{v}_{\text{sp}}-\bm{v}_{\text{fdb}}$ is the velocity error.
2) Attitude Control: We use Hamilton unit quaternions to represent the attitudes. The attitude error is defined as
@@
$\bm{\omega}_{b,\text{sp}}=\text{sign}(e_{Q,w})\cdot\mathbf{K}_{P,Q}\odot\bm{e}_{Q,\text{vec}},$ (12)
@@
$\dot{\bm{\omega}}_{b,\text{sp}}$ $=\mathbf{K}_{P,\omega}\odot\bm{e}_{\omega}+\mathbf{K}_{I,\omega}\odot\int\bm{e}_{\omega}\text{d}t+\mathbf{K}_{D,\omega}\odot\frac{\text{d}\bm{e}_{\omega}}{\text{d}t}$ (13)
$+\mathbf{K}_{F,\omega}\odot\bm{\omega}_{b,\text{ref}}.$
@@
$\mathbf{w}_{b,\text{sp}}:=\begin{bmatrix}\mathbf{f}_{b,\text{sp}}\\ \bm{\tau}_{b,\text{sp}}\end{bmatrix}=\begin{bmatrix}m\mathbf{Q}_{\text{fdb}}^{-1}\otimes(\mathbf{a}_{\text{sp}}-\mathbf{g})\\ \mathbf{J}_{b}\dot{\bm{\omega}}_{b,\text{sp}}+\mathbf{d}_{\text{com}}^{\wedge}\mathbf{f}_{b,\text{sp}}+\bm{\omega}_{b,\text{fdb}}^{\wedge}\mathbf{J}_{b}\bm{\omega}_{b,\text{fdb}}\end{bmatrix},$ (14)
@@
This section presents the details of the proposed 3-stage optimization-based trajectory generation framework based on the above system modeling. The path search and the SFC generation stages are introduced briefly in Section III-A. The trajectory optimization stage is described in detail in Sections III-B to III-D, which is the main contribution of this paper.
@@ III-A Whole-body Safety Constraint
The shape of the vehicle is approximated by a convex polyhedron $\mathcal{P}_{\text{S}}$ that is fixed to $\mathcal{F}_{b}$ and wraps the entire vehicle. The coordinates of its vertices in $\mathcal{F}_{b}$ denoted by $\tilde{\bm{v}}_{l},l=1,\cdots L_{\text{v}}$ are known constants. The safety of the vehicle can be guaranteed as long as the $L_{\text{v}}$ vertices are all in the convex polyhedron $\mathcal{P}$ that represents the safe region. We express $\mathcal{P}$ using linear inequalities:
@@
Our goal is to find the optimal $\mathbf{c}$ and $\mathbf{T}$ that minimize a given objective function, but the high dimensionality of $\mathbf{c}$ may reduce efficiency. By enforcing the optimality conditions in Theorem 2 of [12], the parameters of $\mathbf{z}(t)$ can be transformed from $(\mathbf{c},\mathbf{T})$ to the waypoints $\mathbf{q}:=\begin{bmatrix}\mathbf{q}_{1}&\cdots&\mathbf{q}_{M-1}\end{bmatrix}\in\mathbb{R}^{6\times(M-1)},i=1,\cdots,M-1$ and $\mathbf{T}$. After specifying $(\mathbf{q}$, $\mathbf{T})$, the start condition $\bar{\mathbf{z}}_{o}:=\mathbf{z}^{[s-1]}(t_{0}):=\begin{bmatrix}\mathbf{z}(t_{0})^{\top}&\cdots&\mathbf{z}^{(s-1)}(t_{0})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6s}$, and the end condition $\bar{\mathbf{z}}_{f}:=\mathbf{z}^{[s-1]}(t_{0}):=\begin{bmatrix}\mathbf{z}(t_{M})^{\top}&\cdots&\mathbf{z}^{(s-1)}(t_{M})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6s}$, $\mathbf{c}$ will be determined uniquely in an efficient way by solving the sparse linear system
@@
The first term of the objective function (23a) is the smoothness cost, from now on denoted as $J$, and the second is the time regularization term. Spatial constraints (23d)-(23f) bind the position waypoints to a specific region in the SFC, where $k_{i},i=1,\cdots,M_{\mathcal{P}}$ is the indices of waypoints that should be confined to $\left(\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\right)^{\circ}$ and $11\leq k_{i}<k_{i+1}\leq M-1,i=1,\cdots,M_{\mathcal{P}}-$ 1, as illustrated in Fig. 4. Temporal constraint (23g) ensures that the time allocated to each piece is strictly positive. Inequalities (23h)-(23j) are kinodynamic constraints according to the task requirements and vehicle limits; Safety constraint (23k) confines each trajectory piece to a polyhedron primitive of SFC. We let $\mathcal{P}^{i}:=\{\mathbf{p}\in\mathbb{R}^{3}|\mathbf{n}_{i,k}^{\top}\mathbf{p}-d_{i,k}\leq 0,k=1,\cdots,K_{i}\}$ denote the polyhedron to which the $i$-th piece is assigned.
@@
where $\mathcal{W}_{\star}$ is the weight of the corresponding penalty term. $\kappa\in\mathbb{N}_{+}$ controls the resolution of numerical integration. $\mathcal{V}(\cdot):=\max(\cdot,0)^{3}$ measures the constraint violation on the trajectory at the sampling time $\hat{t}_{ij}:=t_{i-1}+\frac{j}{\kappa}T_{i}$. Problem (24) can be solved by quasi-Newton methods.
III-C Gradient Calculation
The gradients of the objective function w.r.t. the optimization variables are needed to solve the unconstrained optimization problem (24). First, we calculate the gradients w.r.t. $\mathbf{c}$ and $\mathbf{T}$ (refer to Appendix A), Then, the gradients w.r.t. optimization variables $\bm{\xi}$, $\mathbf{q}^{\sigma}$, and $\bm{\tau}$ can be obtained using the formulas derived as [12].
Since the rotation-related quantities such as $\mathbf{R}$ and $\bm{\omega}$ are closely related to the rotation vector $\bm{\sigma}$ defined in Section II-B, the evaluation of these penalty terms and their gradients will vary depending on the rotation parameterization map $\mathbf{R}(\bm{\sigma}):\mathbb{R}^{3}\rightarrow SO(3)$.
III-D Rotation Parameterization
From Section III-B and Appendix A we can find that choosing an appropriate rotation parameterization map $\mathbf{R}(\bm{\sigma})$ is essential for our trajectory generation. There are several commonly used ways to parameterize a rotation in $SO(3)$ as a vector in $\mathbb{R}^{3}$, such as Euler angles and axis-angle (also known as a Lie algebra of $SO(3)$). However, since $\bm{\sigma}$ is a free vector on which we do not apply any hard constraints, it will be ambiguous if $\bm{\sigma}$ represents Euler angles or an axis-angle: two very different $\bm{\sigma}$ values will most likely correspond to the same rotation. Moreover, Euler angle representation has the problem of gimbal lock.
Considering these shortcomings, Euler angle or an axis-angle representation lacks rationality when using polynomial interpolation. In our implementation, we adopt a parameterization method based on Hamilton quaternion representation [14] and stereographic projection, which is more rational as shown in [18]. It uses the homeomorphism between the hyperplane $\mathbb{R}^{3}$ and the hypersphere $\mathbb{S}^{3}$ with one pole removed. A stereographic projection maps an arbitrary vector $\bm{\sigma}\in\mathbb{R}^{3}$ as a unit quaternion $\mathbf{Q}:=\begin{bmatrix}w&x&y&z\end{bmatrix}^{\top}=\begin{bmatrix}w&\mathbf{r}^{\top}\end{bmatrix}^{\top}$ representing the rotation
$\mathbf{R}(\mathbf{Q}):=\begin{bmatrix}1-2y^{2}-2z^{2}&2(xy-wz)&2(xz+wy)\\ 2(xy+wz)&1-2x^{2}-2z^{2}&2(yz-wx)\\ 2(xz-wy)&2(yz+wx)&1-2x^{2}-2y^{2}\end{bmatrix}.$ (25)

### v2

@@ I Introduction
In order to better exploit the potential of OMAVs, a proper trajectory generation framework is indispensable. It needs to generate 6-D $SE(3)$ trajectories and take into account the vehicle’s shape and pose to adapt to cluttered environment. Moreover, the lightweight requirements of MAVs limit the available onboard computing power. So we also expect the designed framework to have lower computational burdens and higher computational efficiency than traditional methods.
@@
This paper proposes an efficient 3-stage 6-D trajectory generation framework for OMAVs in cluttered environments, as shown in Fig. 1. Similar to [11], 3-D polyhedral SFC is used to represent the collision-free region. First, we use sampling-based methods like RRT [14] to search for an initial feasible path from the start to the end point and generate a 3-D SFC based on the initial path. Then, the 6-D trajectory optimization stage efficiently generates a smooth, safe, and dynamically feasible 6-D trajectory connecting the start and the end states with linear time and space complexity. For trajectory representation, we represent a vehicle’s attitude as a Hamilton unit quaternion $\mathbf{Q}\in\mathbb{S}^{3}$ [15] and parameterize it into a free 3-D vector using stereographic projection to eliminate the constraints inherent in the $\mathbb{S}^{3}$ manifold. This 3-D attitude vector is combined with the 3-D position vector to form a 6-D pose vector. The complete $SE(3)$ trajectory is then represented as a piece-wise polynomial of the pose vector over time. The vehicle’s shape and attitude are taken into account to apply whole-body safety constraints to the trajectory. We formulate the trajectory generation problem as a constrained optimization problem and finally transform it into an unconstrained one that can be solved using quasi-Newton methods.
@@
  • Based on 3-D rotation parameterization and the idea of the state-of-the-art trajectory optimization method [12] for traditional underactuated MAVs, we represent rotation trajectories as curves in $\mathbb{R}^{3}$ and present a 6-D rigid body trajectory optimization framework considering whole-body safety and dynamic limits for OMAVs.
  • We achieve efficient trajectory optimization by adopting a stereographic-projection-based rotation parameterization.
  • We propose an efficient 6-D trajectory generation pipeline that leverages the full potential of OMAVs for obstacle avoidance for the first time.
@@ II-B System Modeling of OMAVs
and we select CoM position $\mathbf{p}$, CoM velocity $\mathbf{v}=\dot{\mathbf{p}}$, $\mathcal{F}_{b}$’s orientation $\mathbf{Q}$ (expressed as Hamilton unit quaternion [15]), and $\mathcal{F}_{b}$’s angular velocity $\bm{\omega}$ (all expressed in $\mathcal{F}_{W}$) as its state variables:
$\mathbf{x}:=[\underbrace{p_{x},p_{y},p_{z}}_{\mathbf{p}^{\top}},\underbrace{Q_{w},Q_{x},Q_{y},Q_{z}}_{\mathbf{Q}^{\top}},\underbrace{v_{x},v_{y},v_{z}}_{\mathbf{v}^{\top}},\underbrace{\omega_{x},\omega_{y},\omega_{z}}_{\bm{\omega}^{\top}}]^{\top}.$ (2)
where $\mathbf{x}\in\mathbb{R}^{13}$ is the full state vector. The system’s output is given by the position of its CoM and the orientation of $\mathcal{F}_{b}$ expressed in $\mathcal{F}_{W}$:
$\mathbf{y}(t):=\{\mathbf{p}(t),\mathbf{Q}(t)\}\in SE(3).$ (3)
Expressing a rotation trajectory as a function of a unit quaternion or a rotation matrix w.r.t. time $t$ is intuitive. However, this implies equality constraints inherent in $SO(3)$, which can be troublesome for trajectory optimization. So, we consider parameterizing each rotation $\mathbf{Q}$ as an unconstrained 3-D vector. Denote a rotation trajectory by a curve $\bm{\sigma}(t):[t_{0},t_{M}]\rightarrow\mathbb{R}^{3}$, and the corresponding $\mathbf{Q}(t)$ is determined by a smooth surjection $\mathbf{Q}(\bm{\sigma}):\mathbb{R}^{3}\rightarrow SO(3)$. Then, we express the 6-D trajectory as
@@
where $\mathbf{R}$ is the rotation matrix corresponding to $\mathbf{Q}$; $(\cdot)^{\wedge}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ denotes the skew-symmetric matrix of a 3-D vector and thus $\mathbf{a}\times\mathbf{b}=\mathbf{a}^{\wedge}\mathbf{b},\forall\mathbf{a},\mathbf{b}\in\mathbb{R}^{3}$. The inverse map is denote as $(\cdot)^{\vee}:\mathfrak{so}(3)\rightarrow\mathbb{R}^{3}$. The angular acceleration $\dot{\bm{\omega}}$ can be further calculated according to (5) :
@@
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{Q}\\ \mathbf{v}\\ \bm{\omega}\end{bmatrix}=\Psi_{x}(\mathbf{z},\dot{\mathbf{z}}):=\begin{bmatrix}\mathbf{p}\\ \mathbf{Q}(\bm{\sigma})\\ \dot{\mathbf{p}}\\ \left(\sum_{i=1}^{3}\frac{\partial\mathbf{R}}{\partial\sigma_{i}}\dot{\sigma}_{i}\mathbf{R}^{\top}\right)^{\vee}\end{bmatrix}.$ (7)
@@
Equations (7) and (8) indicate that our selected trajectory representation $\mathbf{z}(t)$ has the properties as a flat output [16]. Therefore, it is convenient to transform constraints on the state $\mathbf{x}\in\mathbb{R}^{13}$ and the input $\mathbf{u}\in\mathbb{R}^{6}$ to constraints on $\mathbf{z}\in\mathbb{R}^{6}$ and $\mathbf{z}$’s finite derivatives using Equ. (7) and Equ. (8).
@@ II-C Position and Attitude Control
$\mathbf{v}_{\text{sp}}=\mathbf{K}_{P,p}\bm{e}_{p},$ (9)
$\mathbf{K}_{P,p}\in\mathbb{R}^{3\times 3}$ is the diagonal gain matrix. The inner velocity control loop uses PID control with a feed-forward design, given as
$\mathbf{a}_{\text{sp}}=\mathbf{K}_{P,v}\bm{e}_{v}+\mathbf{K}_{I,v}\int\bm{e}_{v}\text{d}t+\mathbf{K}_{D,v}\frac{\text{d}\bm{e}_{v}}{\text{d}t}+\mathbf{K}_{F,v}\mathbf{v}_{\text{ref}},$ (10)
where $\mathbf{K}_{P,v},\mathbf{K}_{I,v},\mathbf{K}_{D,v},\mathbf{K}_{F,v}\in\mathbb{R}^{3\times 3}$ are diagonal gain matrices for proportional, integral, derivative, and feed-forward control, respectively. $\bm{e}_{v}:=\bm{v}_{\text{sp}}-\bm{v}_{\text{fdb}}$ is the velocity error.
2) Attitude Control: We use Hamilton unit quaternions to represent the attitudes. The attitude error is defined as the difference between the reference attitude $\mathbf{Q}_{\text{ref}}$ and the attitude feedback $\mathbf{Q}_{\text{fdb}}$
@@
$\bm{\omega}_{b,\text{sp}}=\text{sign}(e_{Q,w})\cdot\mathbf{K}_{P,Q}\bm{e}_{Q,\text{vec}},$ (12)
@@
$\dot{\bm{\omega}}_{b,\text{sp}}$ $=\mathbf{K}_{P,\omega}\bm{e}_{\omega}+\mathbf{K}_{I,\omega}\int\bm{e}_{\omega}\text{d}t+\mathbf{K}_{D,\omega}\frac{\text{d}\bm{e}_{\omega}}{\text{d}t}$ (13)
$+\mathbf{K}_{F,\omega}\bm{\omega}_{b,\text{ref}}.$
@@
$\mathbf{w}_{b,\text{sp}}:=\begin{bmatrix}\mathbf{f}_{b,\text{sp}}\\ \bm{\tau}_{b,\text{sp}}\end{bmatrix}=\begin{bmatrix}m\mathbf{Q}_{\text{fdb}}^{-1}\otimes(\mathbf{a}_{\text{sp}}-\mathbf{g})\otimes\mathbf{Q}_{\text{fdb}}\\ \mathbf{J}_{b}\dot{\bm{\omega}}_{b,\text{sp}}+\mathbf{d}_{\text{com}}^{\wedge}\mathbf{f}_{b,\text{sp}}+\bm{\omega}_{b,\text{fdb}}^{\wedge}\mathbf{J}_{b}\bm{\omega}_{b,\text{fdb}}\end{bmatrix},$ (14)
@@
This section presents the details of the proposed 3-stage optimization-based trajectory generation framework based on the above system modeling. The path search and the SFC generation stages are introduced briefly in Section III-A. The trajectory optimization stage is described in detail in Sections III-B to III-C, which is the main contribution of this paper.
@@ III-A Whole-body Safety Constraint
Similar to [11], the shape of the vehicle is approximated by a convex polyhedron $\mathcal{P}_{\text{S}}$ that is fixed to $\mathcal{F}_{b}$ and wraps the entire vehicle. The coordinates of its vertices in $\mathcal{F}_{b}$ denoted by $\tilde{\bm{v}}_{l},l=1,\cdots L_{\text{v}}$ are known constants. The safety of the vehicle can be guaranteed as long as the $L_{\text{v}}$ vertices are all in the convex polyhedron $\mathcal{P}$ that represents the safe region. We express $\mathcal{P}$ using linear inequalities:
@@
Our goal is to find the optimal $\mathbf{c}$ and $\mathbf{T}$ that minimize a given objective function. We adopt MINCO trajectory representation proposed in [12] to deal with the high dimensionality of $\mathbf{c}$ which may reduce optimization efficiency. By enforcing the optimality conditions in Theorem 2 of [12], the parameters of $\mathbf{z}(t)$ can be transformed from $(\mathbf{c},\mathbf{T})$ to the waypoints $\mathbf{q}:=\begin{bmatrix}\mathbf{q}_{1}&\cdots&\mathbf{q}_{M-1}\end{bmatrix}\in\mathbb{R}^{6\times(M-1)},i=1,\cdots,M-1$ and $\mathbf{T}$. After specifying $(\mathbf{q}$, $\mathbf{T})$, the start condition $\bar{\mathbf{z}}_{o}:=\mathbf{z}^{[s-1]}(t_{0}):=\begin{bmatrix}\mathbf{z}(t_{0})^{\top}&\cdots&\mathbf{z}^{(s-1)}(t_{0})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6s}$, and the end condition $\bar{\mathbf{z}}_{f}:=\mathbf{z}^{[s-1]}(t_{0}):=\begin{bmatrix}\mathbf{z}(t_{M})^{\top}&\cdots&\mathbf{z}^{(s-1)}(t_{M})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6s}$, $\mathbf{c}$ will be determined uniquely in an efficient way by solving the sparse linear system
@@
The first term of the objective function (23a) is the smoothness cost, from now on denoted as $J$, and the second is the time regularization term. Spatial constraints (23d)-(23f) bind the position waypoints to a specific region in the SFC, where $k_{i},i=1,\cdots,M_{\mathcal{P}}$ is the indices of waypoints that should be confined to $\left(\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\right)^{\circ}$ and $11\leq k_{i}<k_{i+1}\leq M-1,i=1,\cdots,M_{\mathcal{P}}-$ 1, as illustrated in Fig. 4. Temporal constraint (23g) ensures that the time allocated to each piece is strictly positive. Inequalities (23h)-(23j) are kinodynamic constraints according to the task requirements and vehicle limits; Safety constraint (23k) confines each trajectory piece to a polyhedron primitive of SFC. We let $\mathcal{P}^{i}:=\{\mathbf{p}\in\mathbb{R}^{3}|\mathbf{n}_{i,k}^{\top}\mathbf{p}-d_{i,k}\leq 0,k=1,\cdots,K_{i}\}$ denote the polyhedron to which the $i$-th piece is assigned. Note that in addition to (23h)-(23k), we can also add more continuous time constraints according to actual needs, such as dealing with the vehicle’s actuator delays or dynamic changes in the environment.
@@
where $\mathcal{W}_{\star}$ is the weight of the corresponding penalty term. $\kappa\in\mathbb{N}_{+}$ controls the resolution of numerical integration. $\mathcal{V}(\cdot):=\max(\cdot,0)^{3}$ measures the constraint violation on the trajectory at the sampling time $\hat{t}_{ij}:=t_{i-1}+\frac{j}{\kappa}T_{i}$. With closed-form gradients, problem (24) can be efficiently solved by quasi-Newton methods.
III-C Rotation Parameterization
From Section III-B we can find that choosing an appropriate rotation parameterization map $\mathbf{R}(\bm{\sigma})$ is essential for our trajectory generation. There are several commonly used ways to parameterize a rotation in $SO(3)$ as a vector in $\mathbb{R}^{3}$, such as Euler angles and axis-angle (also known as a Lie algebra of $SO(3)$). However, since $\bm{\sigma}$ is a free vector on which we do not apply any hard constraints, it will be ambiguous if $\bm{\sigma}$ represents Euler angles or an axis-angle: two very different $\bm{\sigma}$ values will most likely correspond to the same rotation. Moreover, Euler angle representation has the problem of gimbal lock.
Considering these shortcomings, Euler angle or an axis-angle representation lacks rationality when using polynomial interpolation. In our implementation, we adopt a parameterization method based on Hamilton quaternion representation [15] and stereographic projection, which exhibits excellent numerical properties such as accelerated convergence as shown in [19], to make trajectory optimization more efficient. Utilizing the homeomorphism between the hyperplane $\mathbb{R}^{3}$ and the hypersphere $\mathbb{S}^{3}$ with one pole removed, a stereographic projection maps an arbitrary vector $\bm{\sigma}\in\mathbb{R}^{3}$ as a unit quaternion $\mathbf{Q}:=\begin{bmatrix}q_{w}&q_{x}&q_{y}&q_{z}\end{bmatrix}^{\top}=\begin{bmatrix}q_{w}&\mathbf{r}^{\top}\end{bmatrix}^{\top}$ representing the rotation
$\mathbf{R}(\mathbf{Q}):=(q_{w}^{2}-\mathbf{r}^{\top}\mathbf{r})\mathbf{I}+2\mathbf{r}\mathbf{r}^{\top}+2q_{w}\mathbf{r}^{\wedge}$ (25)